In investing $6,600 of a couple's money, a financial planner put some of it into a savings account paying 3% annual interest. The rest was invested in a riskier mini-mall development plan paying 9% annually. The combined interest earned for the first year was $426. How much money was invested at each rate?
Answer by richwmiller(17219)

Total amount of money invested: $6600
x+y=6600,
Total yearly interest for the two accounts is: $426
0.03*x+0.09*y=426
x=6600-y
Substitute for x
0.03*(6600-y)+0.09*y=426
Multiply out
198-0.03*y+0.09*y=426
Combine like terms.
0.06*y=228
Isolate y
y=$ 3800.00 at 9% earns $342 interest
x=6600-y
Calculate x
x=$ 2800.00 at 3% earns $84 interest
Check
0.03*2800+0.09*3800=426
84+342=426
426=426
If this statement is TRUE and neither x nor y is negative then all is well
code int1